Ask
your own question, for FREE!
Mathematics
9 Online
94 x 49 x 3
Still Need Help?
Join the QuestionCove community and study together with friends!
and implies multiplication. What are the chances of getting question 3 incorrectly?
3/16? @bibby
There are 4 possible choices and another 2 all of these or none of these. It's a multiple choice test so ONLY 1 out of all 6 is correct. Thus her chances at get 3 incorrect are 5/6
5/36?
yeah
Still Need Help?
Join the QuestionCove community and study together with friends!
how can i be sure though?
I'm not sure. I just know, because it's basic probability. Logically, it's true. The probability of A and B is always P(a)*P(b) 5/6*1/6=5/36
Just remember these 2 ideas, OR = add, AND = multiply add the probability of getting red, green, and blue
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
9 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal